OpenStudy (anonymous):

HELP!!! Need MAJOR HELP: present value of an annuity of $50, 6%, 6yr, assume n =12.

OpenStudy (anonymous):

@chelsaloo Post the Present Value formula here, then I'll help you :)

OpenStudy (anonymous):

m=P(r/n) 1-(1+r/n) ^nt

OpenStudy (anonymous):

It's one giant fraction looking thing.

OpenStudy (anonymous):

I assume your post asks for Monthly payment?

OpenStudy (anonymous):

First what's r?

OpenStudy (anonymous):

All it says is, "Use a calculator to evaluate the present value of an annuity formula. For the values of the variables m, r, and t. Assume n= 12 (rounding your answer). #50; 6%; 6yr

OpenStudy (anonymous):

Hint: Rate r is something have %

OpenStudy (anonymous):

@chelsaloo which one is r here?

OpenStudy (anonymous):

r=6%

OpenStudy (anonymous):

6% = ...

OpenStudy (anonymous):

-> 6% / 12 = ...?

OpenStudy (anonymous):

is it 1.06?

OpenStudy (anonymous):

never mind. .005?

OpenStudy (anonymous):

Then 1 + r/ 12 = ....?

OpenStudy (anonymous):

1+.06/12=1.005

OpenStudy (anonymous):

72

OpenStudy (anonymous):

=> ( 1 + r/ 12 ) ^ -( 6 * 12) = ....?

OpenStudy (anonymous):

oops.

OpenStudy (anonymous):

12.364

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

1 - [ ( 1 + r/ 12 ) ^ -( 6 * 12) ] = ....?

OpenStudy (anonymous):

Now you've lost me. I don't know what to do with the exponents. It's a negative and I don't remember what to do?!

OpenStudy (anonymous):

No, ( 1 + r/ 12 ) ^ -( 6 * 12) = .6983

OpenStudy (anonymous):

How did you get that??

OpenStudy (anonymous):

Just plug exponent ^ - 72

OpenStudy (anonymous):

Yeah, I know that, but how did you turn it positive??? My cal says it comes out 0 or it can't process it.

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Make sure you plug the negative sign, the one with (-), not the subtraction sign!